# Studying the Impact of Early Test Termination Due to Assertion Failure on Code Coverage and Spectrum-based Fault Localization

### v1

@@ Abstract
Index Terms:
Early Test Termination, Coverage, Fault Localization, Assertion FailureI Introduction
Testing is an effective and practical way of finding bugs and improving software quality. However, due to the scale of modern software and the complexity of tests, it can be difficult for developers to know the cause of test failures and propose fixes. Prior studies have proposed various automated testing techniques that leverage code coverage and test execution information to assist developers with test failure diagnosis or repair. For example, researchers have proposed using code coverage information in failed/passed tests for tasks such as spectrum-based fault localization (SBFL) [1, 2, 3, 4, 5, 6, 7, 8] and automated program repair [9, 10, 11, 12, 13]. For instance, Ochiai [14] and Tarantula [15], SBFL techniques, leverage both passed/failed tests together with their coverage information to identify the faulty location in the program. In modern continuous integration pipelines, SBFL is deployed and triggered automatically when a test fails [16, 17]. Therefore, when assertion failure happens, SBFL provides additional information that helps developers narrow the search space and identify the faulty code effectively.
In testing, assertions are commonly used to verify the program’s expected behavior (e.g., if the returned value of a method equals an expected value). Although it is a recommended practice to use only one assertion in a test to avoid code smell (e.g., Assertion Roulette [18]), prior studies [19, 20] found that many tests may still contain multiple assertions. One issue with tests that have more than one assertion is that when the test fails at an early assertion, the test will terminate at that point and skip the remaining test code. This, in turn, can potentially reduce the code coverage, hindering the effectiveness of SBFL or automated program repair techniques. We refer to the scenario where a test has multiple assertions and fails at an early assertion (i.e., not the last assertion) as early test termination. Understanding the impact of early test termination on test coverage is important for software testing and debugging, particularly for the techniques that rely on coverage information obtained from the testing.
Therefore, we first conduct an empirical study to understand the prevalence of early test termination and examine its impact on the test coverage and effectiveness of SBFL techniques (i.e., Ochiai and Tarantula) that rely on coverage information. We investigated 207 versions collected from six open-source projects used in Defect4J [21] and T-Evos [1]. We formulate the following research questions in our research:
@@
We make our dataset publicly available [22] to encourage future research on this direction.
@@ II-A Early Test Termination
To assist developers with diagnosing the causes of test failures, researchers have proposed various techniques such as SBFL [1, 2, 3, 4, 6, 7, 8]. However, when a test terminates prematurely, the generated test coverage and test failure messages become incomplete, which affects how developers and techniques such as SBFL locate the root cause of failures.
We first formally define early test termination. Suppose there is a test that has statements {, , …, ,…, }. We define as an early terminated test if terminates at a statement , where . A test may terminate early due to various reasons, such as errors in object initialization, method invocation, or assertion. In this paper, we focus the study on early test termination due to assertion failure. Although having multiple assertions in a test can be considered a test smell [18], we observed that this is a common practice in the studied systems (as shown in Table I, 34.6% to 73.3% of the tests have multiple assertions in our studied projects). We also observe that early test termination due to assertion failure happens frequently (see more details in Section IV). In JUnit, if an assertion fails, the test will terminate immediately at that point, skipping the code that follows, and it will report the test as failed. For instance, in Figure LABEL:lst:example1, the test testRootEndpoints has multiple assertions, and it stops at the third assertion. As a result, the remaining test code after the failed assertion would not be executed, which potentially leads to incomplete coverage of the subject under test. Our goal is to study the impact of early test termination due to assertion failure on test coverage and the performance of SBFL. For simplicity, we refer to early test failure due to assertion failure as early test termination and the test where early test termination happens as Testearly in the rest of the paper.
@@ II-B Spectrum-based Fault Localization
In this work, we focus on Spectrum-Based Fault Localization (SBFL) as the subject technique to understand the impact of early test termination. We choose SBFL because it relies directly on the quality and completeness of the collected code coverage, and it is an essential step for automated program repair. SBFL uses statistical formulas to measure the suspiciousness of program units based on the program execution traces. Execution traces, also called program spectra, contain details of the covered statements in failed and passed tests. The suspiciousness scores computed from program spectra are then used to generate a ranked list of program units that are most likely to be responsible for the fault. A number of SBFL techniques have been developed over the decades, such as Tarantula [15], Ochiai [14], DStar [8], and BARINEL [6]. In particular, Ochiai and Tarantula are commonly used SBFL techniques because of their high effectiveness [2, 23]. For instance, Ochiai is calculated as: . The suspiciousness score of a statement is based on the execution of passed and failed tests. Therefore, the performance of SBFL is potentially impacted by early test termination. For instance, suppose a test has two assertions and to test two branches and in the subject under test, respectively. If the test fails at , then only is covered. If no other tests cover , the Ochiai score for all the statements in will be 0. However, may also fail if the test were able to continue executing, and we would miss the fault in because of early test termination.
@@ III-B Data Preparation
For our experiment, we select six open-source projects from benchmark datasets Defects4J [21] and T-Evos [1]. Table I gives an overview of the descriptive statistics of the projects. Defects4J contains real bugs extracted from large open-source projects, which enables controlled experiments in software debugging and testing research [21]. Defects4J has been widely used for evaluating fault localization techniques in the research community [25, 26, 27, 28]. We selected five projects (i.e., jackson, jfreechart, commons-math (math for short), joda-time, and commons-lang (lang for short)) from Defects4J, because they are from different domains and were used as the benchmark to evaluate various SBFL in prior studies [29, 4]. To increase the diversity of our study, we further expanded our dataset with T-Evos, following an existing study [4]. T-Evos provides continuous test execution and failure data to help understand testing practices and evaluate automated testing techniques. We added one more project fastjson from T-Evos to increase the dataset diversity. In T-Evos, fastjson is the one with the highest number of test cases per commit and with the highest lines of test codes. This allows us to study early test termination on projects that have rigorous and mature testing. As some of the projects from T-Evos, such as jackson-core and jfreechart, overlap with Defects4J, we studied these projects using the versions provided by Defects4J. For each subject project, we execute the program on the selected versions provided by the benchmarks. We also verify that the number of tests is sufficient, the program builds successfully, and there is at least one failed test. In total, we collected 207 versions (i.e., commits with failed tests) from six projects for our experiment. Each project has an average number of tests ranging from 396.6 to 4,874. We summarize the basic statistics of our studied projects in Table I (All section).
@@ IV Results of the research questions
A significant portion of test cases have multiple assertions in a single test, ranging from 34.7% to 73.3% in our different studied projects. From Table I, we observe that fastjson has the lowest percentage of tests containing multiple assertions in a single test, at 34.7% on average. In addition, in the two largest projects math and lang, 36.8% and 73.3% of the tests have multiple assertions. Our findings are compatible with prior studies [44, 19], which show that a test having multiple assertions is prevalent in real-world projects and sets the stage for early test termination to occur.
@@ V-C Threats to Validity
Internal Validity In RQ3, we evaluate SBFL using three commonly used metrics (i.e., EXAM, top@k, and MFR). Although there might be other metrics that could be used for our task, these metrics are commonly used in evaluating SBFL techniques [2, 25]. A threat to our study is that the program transformation possibly change the behavior of test cases. To mitigate this threat, we measure the quality of transformed tests, we applied Trycatch and Slicing on passed test cases with multiple assertions and measured their code coverage. Our results show that Original, Trycatch, and Slicing have almost the same code coverage (at most 0.4 lines difference across all tests), which means the behaviour of the tests remains mostly the same. In addition, our transformation does not change the results from passed to failure. See more detailed results in our replication package [22].
@@ VI Related work
Xuan et al. proposed an approach to separate existing test cases into small fractions (called purified test cases) for each assertion, which is similar to Slicing, and to enhance the test oracles to further localize faults [24]. This work is the most closely related to our study. However, different from Xuan et al’s study which focuses on improving the effectiveness of fault localization by purifying test cases, our study focuses on investigating the impact of early test termination. More specifically, we investigated the prevalence of early test termination and provided empirical evidence for it. Moreover, we investigated the impact of early test termination on code coverage and fault localization by comparing it with the baseline in which we eliminated early test termination using Trycatch.
A few studies have been conducted to understand assert practice in testing [45, 44, 46, 19]. Bavota et al. studied 18 systems to analyze the distribution of test smells and show that Assertion Roulette (i.e., a code smell occurs when a test method has multiple non-documented assertions) is presented in 62% of the total JUnit classes [19]. Zamprogno et al. examined more than 30k assertions from 105 projects and also found 17.4% of the test cases have three or more assertions, which is compatible with our finding in RQ1. On the other hand, various solutions were proposed to improve the assert practice in unit testing [47, 48, 49, 50, 51]. For instance, Fraser and Arcuri proposed EvoSuite, which uses mutation testing to produce a set of assertions that maximizes the number of seeded defects in a class that can be revealed by the test cases. Watson et al. employed a Neural Machine Translation (NMT) based approach called Atlas to automatically generate meaningful assert statements providing a test method and a focal method to test the correctness of the focal method [47]. Our study differs from previous ones in that, we concentrate on examining a specific occurrence of test failure, which arises when an assert statement fails in the middle of a test case, thereby preventing the execution of code after that point.
@@ VI-B Investigating the factors that influence testing coverage and the corresponding techniques
Masri et al. investigated the prevalence of factors that impair the effectiveness of coverage-based fault localization [52]. They found that 72% of their studied buggy versions exhibit the condition for failure was met but the program did not fail. More specifically, three conditions must be satisfied for program failure to occur: (1) the defect’s location must be reached, (2) the program’s state must become infected and (3) the infection must propagate to the output. Several studies have been done on this and demonstrate the prevalence of Coincidental correctness and its negative impact on fault detection and localization techniques [53, 54, 55, 52]. For instance, Assi et al. investigated the impact of coincidental correctness on test suite reduction (TSR), test case prioritization (TCP), and SBFL [55]. They found that TSR and TCP, the negative impact of CC was very significant.
@@ VIII Data Availability
We made our dataset publicly available [22] to encourage future research on studying the impact of test quality on fault localization and related approaches.

### v2

@@ Abstract
I Introduction
Testing is an effective and practical way of finding bugs and improving software quality. However, due to the scale of modern software and the complexity of tests, it can be difficult for developers to know the cause of test failures and propose fixes. Prior studies have proposed various automated testing techniques that leverage code coverage and test execution information to assist developers with test failure diagnosis or repair. For example, researchers have proposed using code coverage information in failed/passed tests for tasks such as spectrum-based fault localization (SBFL) [chen2022t, widyasari2022real, wang2011search, chen2022useful, jones2002visualization, abreu2009spectrum, le2015information, wong2016survey] and automated program repair [le2016history, wen2017empirical, lin2017quixbugs, yi2018correlation, liu2021critical]. For instance, Ochiai [abreu2006evaluation] and Tarantula [tarantula], SBFL techniques, leverage both passed/failed tests together with their coverage information to identify the faulty location in the program. In modern continuous integration pipelines, SBFL is deployed and triggered automatically when a test fails [hassan2023uniloc, kochhar2016practitioners]. Therefore, when assertion failure happens, SBFL provides additional information that helps developers narrow the search space and identify the faulty code effectively.
In testing, assertions are commonly used to verify the program’s expected behavior (e.g., if the returned value of a method equals an expected value). Although it is a recommended practice to use only one assertion in a test to avoid code smell (e.g., Assertion Roulette [AssertionRoulette]), prior studies [bavota2012empirical, kim2020empirical] found that many tests may still contain multiple assertions. One issue with tests that have more than one assertion is that when the test fails at an early assertion, the test will terminate at that point and skip the remaining test code. This, in turn, can potentially reduce the code coverage, hindering the effectiveness of SBFL or automated program repair techniques. We refer to the scenario where a test has multiple assertions and fails at an early assertion (i.e., not the last assertion) as early test termination. Understanding the impact of early test termination on test coverage is important for software testing and debugging, particularly for the techniques that rely on coverage information obtained from the testing.
Therefore, we first conduct an empirical study to understand the prevalence of early test termination and examine its impact on the test coverage and effectiveness of SBFL techniques (i.e., Ochiai and Tarantula) that rely on coverage information. We investigated 207 versions collected from six open-source projects used in Defect4J [defects4j] and T-Evos [chen2022t]. We formulate the following research questions in our research:
@@ I Introduction
We make our dataset publicly available [DataRepository] to encourage future research on this direction.
@@ II-A Early Test Termination
To assist developers with diagnosing the causes of test failures, researchers have proposed various techniques such as SBFL [chen2022t, widyasari2022real, wang2011search, chen2022useful, abreu2009spectrum, le2015information, wong2016survey]. However, when a test terminates prematurely, the generated test coverage and test failure messages become incomplete, which affects how developers and techniques such as SBFL locate the root cause of failures.
We first formally define early test termination. Suppose there is a test that has statements {, , …, ,…, }. We define as an early terminated test if terminates at a statement , where . A test may terminate early due to various reasons, such as errors in object initialization, method invocation, or assertion. In this paper, we focus the study on early test termination due to assertion failure. Although having multiple assertions in a test can be considered a test smell [AssertionRoulette], we observed that this is a common practice in the studied systems (as shown in Table I, 34.6% to 73.3% of the tests have multiple assertions in our studied projects). We also observe that early test termination due to assertion failure happens frequently (see more details in Section IV). In JUnit, if an assertion fails, the test will terminate immediately at that point, skipping the code that follows, and it will report the test as failed. For instance, in Figure LABEL:lst:example1, the test testRootEndpoints has multiple assertions, and it stops at the third assertion. As a result, the remaining test code after the failed assertion would not be executed, which potentially leads to incomplete coverage of the subject under test. Our goal is to study the impact of early test termination due to assertion failure on test coverage and the performance of SBFL. For simplicity, we refer to early test failure due to assertion failure as early test termination and the test where early test termination happens as Testearly in the rest of the paper.
@@ II-B Spectrum-based Fault Localization
In this work, we focus on Spectrum-Based Fault Localization (SBFL) as the subject technique to understand the impact of early test termination. We choose SBFL because it relies directly on the quality and completeness of the collected code coverage, and it is an essential step for automated program repair. SBFL uses statistical formulas to measure the suspiciousness of program units based on the program execution traces. Execution traces, also called program spectra, contain details of the covered statements in failed and passed tests. The suspiciousness scores computed from program spectra are then used to generate a ranked list of program units that are most likely to be responsible for the fault. A number of SBFL techniques have been developed over the decades, such as Tarantula [tarantula], Ochiai [abreu2006evaluation], DStar [wong2016survey], and BARINEL [abreu2009spectrum]. In particular, Ochiai and Tarantula are commonly used SBFL techniques because of their high effectiveness [widyasari2022real, lo2010comprehensive]. For instance, Ochiai is calculated as: . The suspiciousness score of a statement is based on the execution of passed and failed tests. Therefore, the performance of SBFL is potentially impacted by early test termination. For instance, suppose a test has two assertions and to test two branches and in the subject under test, respectively. If the test fails at , then only is covered. If no other tests cover , the Ochiai score for all the statements in will be 0. However, may also fail if the test were able to continue executing, and we would miss the fault in because of early test termination.
@@ III-B Data Preparation
For our experiment, we select six open-source projects from benchmark datasets Defects4J [defects4j] and T-Evos [chen2022t]. Table I gives an overview of the descriptive statistics of the projects. Defects4J contains real bugs extracted from large open-source projects, which enables controlled experiments in software debugging and testing research [defects4j]. Defects4J has been widely used for evaluating fault localization techniques in the research community [pearson2017evaluating, shamshiri2015automatically, xuan2016nopol, martinez2016astor]. We selected five projects (i.e., jackson, jfreechart, commons-math (math for short), joda-time, and commons-lang (lang for short)) from Defects4J, because they are from different domains and were used as the benchmark to evaluate various SBFL in prior studies [just2014mutants, chen2022useful]. To increase the diversity of our study, we further expanded our dataset with T-Evos, following an existing study [chen2022useful]. T-Evos provides continuous test execution and failure data to help understand testing practices and evaluate automated testing techniques. We added one more project fastjson from T-Evos to increase the dataset diversity. In T-Evos, fastjson is the one with the highest number of test cases per commit and with the highest lines of test codes. This allows us to study early test termination on projects that have rigorous and mature testing. As some of the projects from T-Evos, such as jackson-core and jfreechart, overlap with Defects4J, we studied these projects using the versions provided by Defects4J. For each subject project, we execute the program on the selected versions provided by the benchmarks. We also verify that the number of tests is sufficient, the program builds successfully, and there is at least one failed test. In total, we collected 207 versions (i.e., commits with failed tests) from six projects for our experiment. Each project has an average number of tests ranging from 396.6 to 4,874. We summarize the basic statistics of our studied projects in Table I (All section).
@@ IV Results of the research questions
A significant portion of test cases have multiple assertions in a single test, ranging from 34.7% to 73.3% in our different studied projects. From Table I, we observe that fastjson has the lowest percentage of tests containing multiple assertions in a single test, at 34.7% on average. In addition, in the two largest projects math and lang, 36.8% and 73.3% of the tests have multiple assertions. Our findings are compatible with prior studies [zamprogno2022dynamic, bavota2012empirical], which show that a test having multiple assertions is prevalent in real-world projects and sets the stage for early test termination to occur.
@@ V-C Threats to Validity
Internal Validity In RQ3, we evaluate SBFL using three commonly used metrics (i.e., EXAM, top@k, and MFR). Although there might be other metrics that could be used for our task, these metrics are commonly used in evaluating SBFL techniques [widyasari2022real, pearson2017evaluating]. A threat to our study is that the program transformation possibly change the behavior of test cases. To mitigate this threat, we measure the quality of transformed tests, we applied Trycatch and Slicing on passed test cases with multiple assertions and measured their code coverage. Our results show that Original, Trycatch, and Slicing have almost the same code coverage (at most 0.4 lines difference across all tests), which means the behaviour of the tests remains mostly the same. In addition, our transformation does not change the results from passed to failure. See more detailed results in our replication package [DataRepository].
@@ VI Related work
Xuan et al. proposed an approach to separate existing test cases into small fractions (called purified test cases) for each assertion, which is similar to Slicing, and to enhance the test oracles to further localize faults [xuan2014test]. This work is the most closely related to our study. However, different from Xuan et al’s study which focuses on improving the effectiveness of fault localization by purifying test cases, our study focuses on investigating the impact of early test termination. More specifically, we investigated the prevalence of early test termination and provided empirical evidence for it. Moreover, we investigated the impact of early test termination on code coverage and fault localization by comparing it with the baseline in which we eliminated early test termination using Trycatch.
A few studies have been conducted to understand assert practice in testing [bai2022assertion, zamprogno2022dynamic, jia2021unit, bavota2012empirical]. Bavota et al. studied 18 systems to analyze the distribution of test smells and show that Assertion Roulette (i.e., a code smell occurs when a test method has multiple non-documented assertions) is presented in 62% of the total JUnit classes [bavota2012empirical]. Zamprogno et al. examined more than 30k assertions from 105 projects and also found 17.4% of the test cases have three or more assertions, which is compatible with our finding in RQ1. On the other hand, various solutions were proposed to improve the assert practice in unit testing [watson2020learning, tufano2022generating, yu2022automated, song2007unitplus, fraser2011evosuite]. For instance, Fraser and Arcuri proposed EvoSuite, which uses mutation testing to produce a set of assertions that maximizes the number of seeded defects in a class that can be revealed by the test cases. Watson et al. employed a Neural Machine Translation (NMT) based approach called Atlas to automatically generate meaningful assert statements providing a test method and a focal method to test the correctness of the focal method [watson2020learning]. Our study differs from previous ones in that, we concentrate on examining a specific occurrence of test failure, which arises when an assert statement fails in the middle of a test case, thereby preventing the execution of code after that point.
@@ VI-B Investigating the factors that influence testing coverage and the corresponding techniques
Masri et al. investigated the prevalence of factors that impair the effectiveness of coverage-based fault localization [masri2009empirical]. They found that 72% of their studied buggy versions exhibit the condition for failure was met but the program did not fail. More specifically, three conditions must be satisfied for program failure to occur: (1) the defect’s location must be reached, (2) the program’s state must become infected and (3) the infection must propagate to the output. Several studies have been done on this and demonstrate the prevalence of Coincidental correctness and its negative impact on fault detection and localization techniques [hierons2006avoiding, wang2009taming, abou2021detrimental, masri2009empirical]. For instance, Assi et al. investigated the impact of coincidental correctness on test suite reduction (TSR), test case prioritization (TCP), and SBFL [abou2021detrimental]. They found that TSR and TCP, the negative impact of CC was very significant.
@@ VIII Data Availability
We made our dataset publicly available [DataRepository] to encourage future research on studying the impact of test quality on fault localization and related approaches.